### v1

# RoboBrain: A Unified Brain Model for Robotic Manipulation from Abstract to Concrete – Supplementary Material –

RoboBrain: A Unified Brain Model for Robotic Manipulation
from Abstract to Concrete
– Supplementary Material –
First Author
Institution1
Institution1 address
firstauthor@i1.org
   Second Author
Institution2
First line of institution2 address
secondauthor@i2.org
Appendix
This supplementary material provides more details of the proposed method and experiment results that are omitted from the manuscript due to the page limit. Sec. A presents additional details of the models and training strategies. Sec. B presents details of training dataset. Sec. C complements more experiment results and analysis. Sec. D shows more visualization results to prove the effectiveness of RoboBrain. Sec. E introduces more details about the construction of ShareRobot dataset.
Appendix A Details of Models and Training
Model Setting. RoboBrain is built upon the LLaVA [LLaVa] framework and consists of three main components: the visual encoder, projector, and large language model (LLM).
For the visual encoder, we utilized the SigLIP [siglip] model, specifically the siglip-so400m-patch14-384, which is pre-trained on WebLi [pali] at a resolution of 384x384. The SigLIP model improves upon traditional CLIP [clip] architectures by employing a sigmoid loss function that operates solely on image-text pairs, eliminating the need for global normalization of pairwise similarities. This enhancement allows for more efficient scaling of batch sizes while maintaining performance, even at smaller scales. SigLIP has 27 hidden layers and processes input images using patches of size 14x14, resulting in 729 visual tokens per image. The projector consists of a two-layer MLP [llavanext] that projects the visual tokens obtained from the visual encoder to the dimensions of the text embeddings. For the LLM, we adopted the Qwen2.5-7B-Instruct [qwen2_5] model, which is a state-of-the-art open-source LLM that is part of the latest Qwen series [Qwen]. It features 28 hidden layers and supports long-context inputs of up to 128K tokens, providing multilingual capabilities across more than 29 languages.
In Stage 4, we introduced LoRA [lora] to train RoboBrain, enabling it to acquire affordance perception and trajectory prediction capabilities. LoRA is a technique that allows for parameter-efficient fine-tuning of large models by adding low-rank parameter matrices to existing layers. We incorporated LoRA modules with a rank of 64 into the feed-forward network layers of both the Projector and the LLM, freezing all parameters except those of the LoRA modules during training.
Training Setting. In the main text of the paper, we employed a staged training strategy, with complete settings presented in Tab. 1. We primarily referenced the training strategy of LLaVA-Onevision [li2024llava], a state-of-the-art multimodal large language model, and built upon this foundation to expand the robotic training phase. During the entire training phase, we conducted all experiments on a cluster of servers, each equipped with 8$\times$×A800 GPUs.
Table 1: Detailed configuration for each training stage of the RoboBrain.
Stage-1 Stage-1.5 Stage-2 Stage-3 Stage-4
Single-Image OneVision A-LoRA T-LoRA
[…]
#Tokens 729 Max 729$\times$5 Max 729$\times$37 Max 729$\times$37 Max 729$\times$37 Max 729$\times$37 Max 729$\times$37
Model Trainable Projector Full Model Full Model Full Model Full Model A-LoRA T-LoRA
#Tunable Parameters 17.0M 8.0B 8.0B 8.0B 8.0B 28.0M 28.0M
Training Per-device Batch Size 8 2 1 1 1 4 4
Gradient Accumulation 1 2 2 2 2 2 2
[…]
Epoch 1 1 1 1 1 1 1
Optimizer AdamW AdamW AdamW AdamW AdamW AdamW AdamW
Deepspeed Zero3 Zero3 Zero3 Zero3 Zero3 Zero2 Zero2
Weight Decay 0 0 0 0 0 0 0
Warmup Ratio 0.03 0.03 0.03 0.03 0.03 0.03 0.03
LR Schedule cosine cosine cosine cosine cosine cosine cosine
Projector Type mlp2x_gelu mlp2x_gelu mlp2x_gelu mlp2x_gelu mlp2x_gelu mlp2x_gelu mlp2x_gelu
Vision Select Layer -2 -2 -2 -2 -2 -2 -2
Patch Merge Type spatial_unpad spatial_unpad spatial_unpad spatial_unpad spatial_unpad spatial_unpad spatial_unpad
Frames Upbound - - - 32 32 32 32
Max Seq Length 8192 32768 32768 32768 32768 4096 4096
GPU Nums 16*8 16*8 20*8 20*8 22*8 4*8 4*8
Appendix B Details of Training Dataset
In the main body of the paper, we emphasize the importance of the training data and the proportion of robotic data. In this section, we will provide a detailed overview of the training data and its sources. The distribution of the entire training dataset is illustrated in Fig. 1.
Figure 1: The distribution of the entire training dataset.
  • LCS-558K is a subset of the LAION/CC/SBU dataset [cc3m, laion], specifically designed as the LLaVA Visual Instruct Pretrain [LLaVa] Dataset. This dataset has been carefully filtered to achieve a more balanced distribution of concept coverage, ensuring diverse and representative visual content. The primary purpose of LCS-558K is to facilitate the alignment between the visual encoder and the LLM, enabling the LLM to comprehend visual information.
  • Image-4M comprises 8 data sources, including 3 from the LLaVA-Recap series [li2024llava-next]: BLIP558K, COCO118K, and CC3M, as well as UReader [ye2023ureader], Instruct Azure DC [li2024llava-next], Evol-Instruct [chen2024allava], and SynthDog [kim2022donut] We utilized the download links provided by the LLaVA-OneVision team for the data acquisition.
  • SI-3.2M [^1] Due to the unavailability of certain datasets, the actual data used amounts to 3.1M. [li2024llava] consists of 3.2 million samples, carefully curated to support multimodal learning. It includes subsets from existing datasets such as Cambrian [tong2024cambrian], Cauldron [li2024llava-next], and UReader [ye2023ureader], which were subjected to cleaning and re-annotation to ensure data quality. Additionally, it incorporates single-image data from sources like AI2D [ai2d] and OKVQA [marino2019okvqa], alongside a newly compiled single-image collection designed to achieve a balanced and diverse dataset.
  • OV-1.6M [^2] Due to the vague descriptions and missing key information regarding dataset filtering in the original paper, we ended up using 2.4M data. [li2024llava] comprises 1.6 million samples, which includes approximately 800K high-quality samples re-sampled from earlier SI-3.2M datasets with a data replay strategy, ensuring improved data reliability and relevance. Additionally, the dataset incorporates M4-Instruct data to enrich instructional learning tasks. A significant component of OV-1.6M is its video data, which has been released alongside LLaVA-video data. The video subset used in the dataset is specifically aligned with the previous annotation format, providing a diverse multimodal resource for advancing vision-language learning.
  • RoboVQA-800K [RoboVQA] consists of realistic data gathered from various user requests, utilizing different embodiments including robots, humans, and humans equipped with grasping tools. The dataset features 5,246 long-horizon episodes and 92,948 medium-horizon episodes of robotic tasks, with each episode accompanied by corresponding image and text prompt inputs. The primary purpose of RoboVQA-800K is to enhance RoboBrain’s reasoning capabilities in robotic-related scenarios.
  • ScanView-318K totals 318K samples, which integrates data from several high-quality sources, including MMScan-224K [mmscan], 3RScan-43K [3rscan], ScanQA-25K [scanqa], and SQA3D-26K [sqa3d], each contributing unique strengths. MMScan-224K provides multimodal scene data with detailed annotations, such as object segmentation and textual descriptions. 3RScan-43K offers 3D reconstructions and semantic annotation. ScanQA-25K includes question-answer pairs based on 3D scanned environments. SQA3D-26K focuses on spatial question answering. Together, these datasets provide diverse scene-scanning image data, long video sequences, and high-resolution samples, equipping models with fine-grained environmental perception and reasoning abilities.
Appendix C Complementary Experiments
In this section, we present the complete experiments and results that are omitted from the manuscript due to page limitations. This includes an exploration of the impact of incorporating ShareRobot on training, the effects of varying proportions of robotic data in the training dataset, and more comprehensive results comparing RoboBrain with the baselines on both general and robotic benchmarks.
Additionally, we explore the impact of different architectures and pre-trained VLMs, as well as different LLM backbones on our experimental results. We also conduct ablation studies at various stages to meticulously analyze the contributions of each stage to overall performance.
Table 2: Performance comparison on multiple general benchmarks.
Dataset Split RoboBrain (Ours) GPT-4V [gpt4v] LLaVA-OV-7B [llavanext] InternVL2-8B [internvl] Qwen2-VL-7B [qwen2vl] GPT-4o [gpt4o]
A12D[ai2d] test 82.03 78.2 81.4 83.8 - 94.2
ChartQA[chartqa] test 80.48 78.5 80 83.3 83 85.7
DocVQA[docvqa] test 88 88.4 87.5 91.6 94.5 92.8
TextVQA[textvqa] val 75.85 - 71.07 77.4 84.3 -
MMMU[mmmu] val 49 56.8 48.8 51.8 54.1 69.1
MMStar[mmstar] test 61.23 57.1 61.7 61.5 60.7 63.9
OCRBench[ocrbench] - 677 656 697 794 845 805
RealWorldQA[realworldqa] test 68.89 61.4 66.3 64.4 70.1 58.6
SeedBench[seedbench] image 71.03 49.9 75.4 76.2 - 76.2
MMbench[mmbench] en-dev 81.52 81.3 83.2 - - 83.4
MMbench[mmbench] en-test 80.44 75 80.8 81.7 83 -
MME[mme] test 2084 1926 1998 2210 2327 -
C.1 More Results on General Benchmarks
To evaluate performance on general tasks in real-world scenarios, as is commonly done with MLLMs [gpt4v, llavanext, internvl, qwen2vl, gpt4o], we conducted experiments using a diverse set of image benchmarks summarized in Table 2. We leveraged the comprehensive evaluation toolkit, LMMs-Eval[lmms, zhang2024unveiling], to evaluate RoboBrain’s performance on general benchmarks. These benchmarks are categorized into three classes:
  • Chart, Diagram, and Document Understanding. As key visual formats for structured OCR data, benchmarks such as AI2D [ai2d], ChartQA [chartqa], DocVQA [docvqa], and OCRBench [ocrbench] were utilized. Open-source models like InternVL2-8B [internvl] and LLAVA-OV-7B [llavanext] have demonstrated comparable performance to closed-source models such as GPT-4V [gpt4v]. For RoboBrain, despite being optimized primarily for multidimensional robotic tasks, it surpasses LLAVA-OV-7B [llavanext] and GPT-4V [gpt4v] on these benchmarks, achieving a significant improvement in structured OCR tasks, with the only exceptions being DocVQA [docvqa], where it performs slightly lower than GPT-4V [gpt4v], and OCRBench [ocrbench], where it falls slightly behind LLAVA-OV-7B [llavanext].
  • Visual Perception and Multi-domain Reasoning. This category focuses on complex visual perception and multidisciplinary reasoning tasks. Benchmarks for visual perception include MMStar [mmstar], MMBench [mmbench], and MME [mme], while reasoning benchmarks include MMMU [mmmu] and SeedBench [seedbench]. RoboBrain demonstrates comparable performance to GPT-4V [gpt4v] and LLAVA-OV-7B [llavanext] across multiple benchmarks.
  • Real-world Understanding and Interaction. Evaluating MLLMs [gpt4v, llavanext, internvl, qwen2vl, gpt4o] as general-purpose assistants in real-world settings is crucial, as these scenarios extend beyond controlled environments. For this, the RealworldQA [realworldqa] benchmark was utilized. Results indicate that RoboBrain not only outperforms open-source models like LLAVA-OV-7B [llavanext] and InternVL2-8B [internvl], but also exceeds closed-source models such as GPT-4V [gpt4v] and GPT-4o [gpt4o], showcasing its extensive knowledge base and strong generalization capabilities.
Table 3: Performance comparison on RoboVQA, OpenEQA and ShareRobot Benchmarks.
Dataset Split / Metric RoboBrain (Ours) GPT-4V [gpt4v] LLaVA-OV-7B [llavanext] RoboMamba [Robomamba] Qwen2-VL-7B [qwen2vl]
RoboVQA[RoboVQA] BLEU1 72.05 32.23 38.12 54.9 33.22
BLEU2 65.35 26.51 33.56 44.2 26.11
BLEU3 59.39 24.65 31.76 39.5 20.98
BLEU4 55.05 23.94 30.97 36.3 17.37
OpenEQA[OpenEQA] OBJECT-STATE-RECOGNITION 70.4 63.2 72.02 - 72.06
OBJECT-RECOGNITION 49.54 43.4 51.73 - 61.91
FUNCTIONAL-REASONING 57.14 57.4 55.53 - 54.23
SPATIAL-UNDERSTANDING 46.46 33.6 48.98 - 50.39
ATTRIBUTE-RECOGNITION 66.7 57.2 75.52 - 73.88
WORLD-KNOWLEDGE 53.12 50.7 56.46 - 57.3
OBJECT-LOCALIZATION 47.45 42 45.25 - 47.29
ShareRobot (Eval) DISCRIMINATIVE 99.02 - 57.9 - 76.47
FUTURE-PREDICTION 72.92 - 13.1 - 8.04
GENERATIVE 32.43 - 5.44 - 4.63
PAST-DESCRIPTION 37.07 - 4.4 - 13.65
PLANNING-REMAINING 71.29 - 24.5 - 7.56
PLANNING-TASK 52.43 - 25 - 36.34
PLANNING-WITH 91.95 - 44.25 - 45.12
SUCCESS 61.7 - 58.5 - 54.63
C.2 More Results on Robotic Benchmarks.
To evaluate RoboBrain’s performance on robotic capabilities in real-world scenarios, we selected RoboVQA [RoboVQA], OpenEQA [OpenEQA], and the test set of ShareRobot, extracted from the proposed ShareRobot dataset, as robotic benchmarks for multi-dimensional assessment, as shown in Table 3. The chosen baselines include MLLMs such as GPT-4V [gpt4v], LLaVA-OV-7B [llavanext], and Qwen2-VL-7B [qwen2vl], as well as robotic models like RoboMamba [Robomamba]. Detailed descriptions of the three selected robotic benchmarks and the analysis of each results are provided below:
[…]
  • OpenEQA [OpenEQA] provides a robotics VQA benchmark with over 1,600 high-quality human-generated questions drawn from more than 180 real-world scenes, targeting the task of Embodied Question Answering (EQA) for environment understanding. For fairness, we evaluated all models using the prompt templates and the LLM-Score metric provided by OpenEQA [OpenEQA]. Based on the evaluation results, our proposed model, RoboBrain, outperforms GPT-4V [gpt4v] overall and achieves comparable performance to other baselines. In the future, we plan to further enhance RoboBrain’s spatial intelligence to improve its generalization across scenes.
  • ShareRobot (Eval) provides a cross-scene and cross-embodiment robotics benchmark consisting of 2,050 VQA pairs, drawn from 102 diverse scenes (e.g., bedroom, laboratory, kitchen, office) and covering 12 different robot bodies. Similar to RoboVQA [RoboVQA], we categorized various robotic VQA tasks into planning, planning with context, planning remaining steps, future prediction, generative affordance, past description, success (positive/negative), and discriminative affordance (positive/negative). Unlike RoboVQA benchmark [RoboVQA], we utilized GPT-4o [gpt4o] to score the evaluation results instead of BLEU metrics for each task, aiming for more accurate performance assessment. Based on the results, our proposed model, RoboBrain, outperforms all baselines, demonstrating its exceptional planning capabilities across diverse scenes and embodiments.
C.3 Effectiveness of ShareRobot
In this subsection, we investigate the effectiveness of the proposed ShareRobot dataset for training RoboBrain. We maintain the ratio of robotic data to general data used in the main body of the paper, approximately 4:6. Based on the original data source proportions, we randomly sampled 200K samples, which include:
  • Exp A consists of 40% robotic data, with 20% sourced from ShareRobot and 20% from other robotic sources, along with 60% general data.
  • Exp B consists of 40% robotic data, excluding ShareRobot, with the same other robotic data resampled as in Experiment A, resulting in a total of 40%. It also includes 60% general data, which is identical to that of Exp A.
We conducted a complete epoch for all the experiments mentioned above. The results are presented in Tab 4. As shown in the table, the inclusion of ShareRobot data enhances the model’s performance compared to scenarios without ShareRobot.
Table 4: Experimental results for effectiveness of ShareRobot and different robot data proportion.
Exp. Name General Data (%) Robotic Data (%) General Benchmarks Robotic Benchmarks Average
OneVision ShareRobot Others Type-1 Type-2 Type-3 RoboVQA[RoboVQA] OpenEQA[OpenEQA] ShareRobot
EXP A 60% 20% 20% 62.44 71.98 70.33 48.29 58.74 63.11 62.48
EXP B 60% 0% 40% 62.36 71.38 66.01 49.20 57.96 27.03 55.66
EXP C 70% 15% 15% 62.73 72.19 68.10 45.96 56.59 61.73 61.22
EXP D 60% 20% 20% 62.44 71.98 70.33 48.29 58.74 63.11 62.48
EXP E 50% 25% 25% 62.28 71.25 66.54 49.34 58.76 63.35 61.92
EXP F 40% 30% 30% 62.39 71.61 68.37 49.22 56.24 64.57 62.07
EXP G 30% 35% 35% 62.69 71.92 69.54 47.74 55.72 65.22 62.14
C.4 Effectiveness of Robot Data Proportion
In this subsection, we investigate the effectiveness of the ratio of robotic data (including ShareRobot) to general data used in training RoboBrain. We maintain a constant total training dataset size of 200K while varying the sampling proportions of robotic and general data. The configurations are as follows:
  • Exp C utilizes a ratio of 3:7, comprising 30% robotic data and 70% general data.
  • Exp D utilizes a ratio of 4:6, comprising 40% robotic data and 60% general data, same to Exp A.
  • Exp E utilizes a ratio of 5:5, with 50% robotic data and 50% general data.
  • Exp F utilizes a ratio of 6:4, featuring 60% robotic data and 40% general data.
  • Exp G utilizes a ratio of 7:3, containing 70% robotic data and 30% general data.
We conducted a complete epoch for all the experiments mentioned above. The results are presented in Tab 4. As shown in the table, a 4:6 ratio of robotic data represents a good choice for training data, effectively balancing performance on both the robotic and the general benchmark.
Table 5: Additional Experimental Results. “SFT Data (G:R)” indicates the ratio of training data for fine-tuning VLMs, where “G” represents general VQA data and “R” denotes robot data (with half being ShareRobot). The total dataset size is 1.47M.
[…]
C.5 Different Architecture and Pre-trained VLMs
To validate the effectiveness of different architecture and pre-trained VLMs and training data in the stage 3 training setup, we selected LLaVA-OneVision-7B [li2024llava], OpenVLA-7B [OpenVLA], and Qwen2VL-7B [qwen2vl], each representing a distinct architecture among VLMs, and conducted SFT using the same proportion of training data described in the main text. As shown in Tab. LABEL:tab_additional (a), the results demonstrated that incorporating ShareRobot can significant performance improvements. For unaligned VLMs like LLaVA 1.5 [llava-1.5] and OpenVLA, we first aligned the MLP using BLIP-558k [li2024llava-next]; other models were directly finetuned.
C.6 Different LLM Backbones
To demonstrate the effectiveness of different LLM backbones when finetuned on the ShareRobot dataset, we conducted experiments using four distinct LLMs [Qwen, llama, vicuna, mistral]. These models were finetuned using the ShareRobot data, and the experimental results are summarized in Tab. LABEL:tab_additional (b). The findings indicate that different LLMs benefit from the ShareRobot data.
C.7 Ablation Studies of Different Stages
We present the evaluation results for each stage in Tab. LABEL:eval. The results demonstrate that staged training from stage 1 to stage 3 consistently and effectively improves the model’s planning performance, while stage 4 enhances the model’s affordance and trajectory capabilities.
Table 6: Additional Evaluation Results.
[…]
Appendix D More Qualitative Results
In this section, we provide additional visual results for planning, affordance perception, and trajectory prediction. This includes the presentation of both positive and negative samples, as well as further analysis.
D.1 Visualization on Planning
Here, we present additional embodied planning for robotic tasks generated by RoboBrain, as shown in Fig. 2. In this figure, we demonstrate the planning results of RoboBrain for four distinct robotic manipulation tasks: ”Water plants”, ”Put the pot in the drawer”, ”Cluster blocks of the same color into different corners”, and ”Clean the desk”, where the first three are categorized as good cases, and the last one as a bad case. Additionally, the model provides a rationale and detailed explanation for each step of the planning process across all four cases.
From the first three planning cases, it is evident that RoboBrain effectively utilizes environmental information and the states of interactive objects—captured from first- or third-person perspective images—to generate task plans for various types of robotic manipulation tasks. Notably, in the ”Cluster blocks of the same color into different corners” task, RoboBrain not only analyzes the number of blocks of each color on the table in Steps 1 and 2 but also provides detailed sub-steps in Step 3, i.e., ”Move the objects to form clusters”. Specifically, it plans the movement of blocks of four different colors to their designated locations: ”top left corner”, ”top right corner”, ”bottom left corner”, and ”bottom right corner”. The exceptional task generalization capability of RoboBrain in planning further validates the effectiveness of our training dataset—including the proposed ShareRobot dataset—and the Multi-Phase training strategy.
We also present a bad case for RoboBrain, namely the ”Clean the desk” task. In this case, the first-person perspective image depicts a work desk spilled with coffee, where the main objects of focus include a ”tissue box”, a ”tipped-over coffee cup”, and the ”spilled coffee liquid”. The errors in the planning results inferred by RoboBrain are summarized as follows: (1) Object recognition error. The only available object for wiping the desk in the image is a ”tissue”, rather than a ”disinfectant wipe”. (2) Omission of critical steps. Before wiping the desk, it is necessary to extract a tissue from the tissue box. However, this step is missing in RoboBrain’s planning. (3) Action decision deviation. In Step 2, i.e., ”Wipe down the desk with a disinfectant wipe”, the detailed description states, ”Start from one end of the desk and move to the other”. This implies that RoboBrain fails to prioritize wiping the ”spilled coffee liquid” specifically, focusing instead on cleaning ”the entire desk”. The primary cause might be the similarity in color between the desk and the spilled coffee, making it difficult for the model to distinguish.
In our extensive testing, although a small number of unreasonable bad cases like the one described above were observed, RoboBrain demonstrated robust planning capabilities in the vast majority of cases. This provides a solid foundation for executing long-horizon manipulation tasks.
Figure 2: Additional embodied planning of RoboBrain. (a)$\sim$(c) show some good cases of RoboBrain’s embodied planning, while (d) shows its bad case. More detailed analysis can be found in Sec.D.1.
D.2 Visualization on Affordance
Here, we present the visualizations of RoboBrain’s perception of affordance areas, as shown in Fig.3. The text below each subfigure indicates the task instructions, while the red bounding boxes represent the affordance areas predicted by the RoboBrain model. The visualizations in the first three rows demonstrate that our RoboBrain model can effectively provide reasonable affordance areas based on human instructions and visual information. For example, given the instruction “drink_with the bottle”, RoboBrain can determine that the bottle cap is in a closed state, thus providing affordance information for the cap area. This highlights RoboBrain’s strong understanding of abstract instructions.
We also present several failure cases, as illustrated in the fourth row of Fig.3. These include misidentified objects, interference from other objects in the scene, and instances where no objects were recognized. These issues may stem from the model’s limited ability to perceive and localize in noisy environments.
Figure 3: Additional visualizations of diverse affordance areas. The text below each subfigure indicates the task instructions, while the red bounding boxes represent the affordance areas predicted by the RoboBrain model. The visualizations in the first three rows demonstrate that our RoboBrain model effectively identifies reasonable affordance areas based on human instructions and visual information. The fourth row presents several failure cases, which may stem from the model’s lack of ability to perceive and localize in noisy environments. This limitation could be attributed to the absence of such scenarios in the training data used during Stage 4. The complete prompt provided to RoboBrain is: ”You are a Franka robot using joint control. The task is \$TASK. Please predict all possible affordance areas of the end effector.” Here, \$TASK represents specific task instructions, such as ”drink with the cup.”
D.3 Visualization on Trajectory
Here, we present additional visualizations generated by RoboBrain using start points, as shown in Fig.4. In this figure, the red-to-purple gradient curves represent the ground truth, while the green-to-blue gradient curves indicate the predicted trajectories. For clarity, waypoints are omitted. The first three rows demonstrate that, regardless of the complexity of the end-effector trajectory, RoboBrain accurately predicts 2D trajectories based on visual observations and task instructions. These predictions closely align with the structure of the ground truth and remain executable.
Additionally, RoboBrain’s predictions often capture the essential features of the trajectories, leading to smoother and potentially more efficient paths compared to the ground truth. This improvement may stem from the inherent variability in the robot’s actual trajectories, which can include redundant waypoints under similar manipulation scenarios. By learning from a large, embodied dataset and utilizing the reasoning capabilities of large language models, RoboBrain is able to infer effective and optimized execution paths.
The visualizations in the third row further suggest that RoboBrain avoids overfitting; it generalizes well across different scenarios, producing trajectories that are both executable and reasonable.
We also present several failure cases, as shown in the fourth row of Fig. 4. These include the robot’s end-effector failing to accurately locate the cup, neglecting the articulated nature of the fridge door while opening it, and not accounting for the deformable properties of clothing during folding. These examples highlight the need for improved spatial perception, as well as the incorporation of object-specific physical constraints and world knowledge to generate more feasible and realistic trajectories.
Figure 4: Additional visualizations of diverse 2D trajectories. The red-to-purple gradient curves represent the ground truth, while the green-to-blue gradient curves indicate the predicted trajectories. The visualizations in the first two rows demonstrate that our RoboBrain model effectively generates end-effector manipulation curves based on the robot’s observations and task instructions. The third row shows that RoboBrain is not merely fitting trajectories but also exhibits the ability to generate more reasonable and feasible curves. The fourth row presents some failure cases, which stem from a lack of spatial awareness and world knowledge. These limitations result in an inability to accurately localize the objects involved in interactions, account for physical constraints, and adapt to the variability of deformable objects.
Appendix E Details of ShareRobot Dataset
In the previous section, we introduced the process of collecting and annotating our ShareRobot dataset. Here, we will provide detailed prompts for data labeling and display examples.
E.1 Prompts
The prompts used for data planning labeling are shown in Fig.5.
Figure 5: Additonal visualizations of prompts for Gemini. The prompts encapsulate the task description for robotic arm action recognition, the components of the target, and the desired response format. Additionally, an example is included to assist Gemini in understanding the specific task.
E.2 High-level Descriptions Examples
In our ShareRobot dataset, there are 10,290 long-horizon high-level descriptions. We provide the 50 most frequently occurring ones below.
  • Closing a drawer
  • Opening a drawer
  • Opening a cabinet door
  • Dragging a strainer across a table
  • Picking up a bowl
  • Inserting a three-pronged object into its matching slot
  • Inserting a double-square object into its matching slot
  • Opening a door
  • Closing a cabinet door
  • Inserting a star-shaped object into its corresponding slot
  • Opening a laptop
  • Inserting an oval object into its corresponding slot
  • Picking up a ketchup bottle from a table
  • Moving a banana from a plate to a table
  • Closing a door
  • Switching a light switch
  • Inserting an arch-shaped object into its corresponding slot
  • Inserting a square-circle object into its matching slot
  • Dragging a strainer backwards
  • Dragging a mug from left to right
  • Dragging a mug forward
  • Picking up a red object from a table
  • Placing a ketchup bottle onto a plate
  • Placing a bowl inside an oven
  • Inserting a hexagonal object into its corresponding slot
  • Closing a microwave door
  • Moving a banana from a table to a plate
  • Turning on a toaster
  • Opening a microwave
  • Closing an oven door
  • Making tea
  • Dragging a strainer forward
  • Placing a bowl into an oven
  • Picking up a banana and placing it in a mug
  • Inserting an arch-shaped object into its matching slot
  • Closing a tea container
  • Inserting a green object into a designated slot
  • Picking up a banana and placing it in a strainer
  • Moving a cloth to the left side of a table
  • Dragging a mug backwards
  • Placing a bottle into a pot
  • Dragging a strainer forwards
  • Inserting a rectangular prism into its matching slot
  • Opening a refrigerator
  • Opening a tea container
  • Opening a double door
  • Inserting a cylinder into a matching hole
  • Picking up a piece of toast
  • Dragging a mug from left to right on a table
  • Closing a refrigerator door
E.3 Low-level Instructions Examples
Our ShareRobot dataset contains 28,181 low-level instructions. The top 50 frequency occurrences are displayed below.
  • Grasp the ketchup bottle
  • Reach for the ketchup bottle
  • Grasp the banana
  • Lift the ketchup bottle
  • Lift the banana
  • Reach for the strainer
  • Reach for the banana
  • Reach for the mug
  • Grasp the mug
  • Lift the pot
  • Lift the bowl
  • Pull the drawer open
  • Reach for the bowl
  • Reach for the pot
  • Grasp the strainer
  • Reach for the drawer handle
  • Grasp the handle
  • Lift the spoon
  • Grasp the bowl
  • Reach for the spoon
  • Place the ketchup bottle on the table
  • Release the banana
  • Reach the drawer
  • Place the banana on the table
  • Lift the mug
  • Reach the cabinet door
  • Grasp the pot
  • Grasp the strainer
  • Grasp the drawer handle
  • Release the mug
  • Grasp the pot
  • Grasp the spoon
  • Place the mug down
  • Move the banana towards the table
  • Grasp the bowl
  • Pull the drawer closed
  • Move towards the bowl
  • Reach for the cloth
  • Release the pot
  • Grasp the bottle
  • Lift the cloth
  • Lift the red object
  • Grasp the banana
  • Lift the butter
  • Reach for the banana on the table
  • Place the ketchup bottle on the plate
  • Grasp the drawer handle
  • Move the ketchup bottle towards the plate
  • Move towards the red object
  • Place the banana on the plate